A ferry leaves a harbor and travels to a resort island at an average speed of 22 mph. On the return trip, the ferry travels at an average speed of 10 mph because of fog. The total time for the trip is 4 h. How far is the island from the harbor?
Found 2 solutions by richwmiller, drk:
Answer by richwmiller(17219)

22*x=10(4-x)
22x=40-10x
32x=40
x=40/32
x=5/4=1.25
22*1.25=10*2.75
27.5=27.5
27.5 miles from the harbor

Answer by drk(1908)

This is an RTD problem. Here is a table based on the given information
location . . . . . rate . . . . .time . . . . . distance
resort . . . . . . . .22 . . . . .t. . . . . . . . . . 22t
harbor . . . . . . .10 . . . . .4-t. . . . . . . . . 40-10t
total. . . . . . . . . . . . . . . . 4
since distances are =, set
22t+=+40-10t
add 10t to get
32t+=+40
divide by 32 to get
t+=+1.25 hours
we want distance, so
22*1.25 = 27.5 miles